# The occurrence of small, short-period planets younger than 200 Myr with TESS

### 05 Mar 2024 v1

@@ 6.1 Forward Modeling
We perform a forward modeling test to compare the demographic of planets around young stars against that of the Kepler distribution. We synthesize a population of planets using the Kepler statistics as presented in Kunimoto & Matthews (2020) and compute the expected number of recovered planets as per our survey efficiency. The forward modeling process allows us to compare the expected characteristics of the planet population against those that we identify from the young stellar sample.
We adopt the Kunimoto & Matthews (2020) Kepler occurrence rates for F, G, and K stars. We draw the occurrence rates assuming half Gaussians based on the asymmetric uncertainties in Kunimoto & Matthews (2020) for each of the simulations. For each star in our parent sample, we simulate a planetary system based on its Kepler period-radius planet frequency grid for its specific stellar type.444The definition of the stellar type boundaries are adopted to be the same as Kunimoto & Matthews (2020) For the M-dwarfs in our sample, we use the occurrence rates for K stars to approximate their planet occurrence rate.. The planets are first randomly assigned into a period-radius cell based on the 2-dimensional cumulative probability. Then, their period and radius are drawn within the cell assuming log-uniform distributions in both axes. We then account for transit probability, and detection probability as per the injection and recovery exercise described in Section 5. For each planet, the transit probability is calculated as , as we reject near-grazing candidates with impact parameters of due to their difficulty for confirmation. Detection probability is computed as the percentage of recovered planets within a single period-radius bin as shown in Figure 6 from our injection and recovery exercise. This forward modeling process is performed 100 times each for the QLP and SPOC samples to quantify its associated uncertainties. The resulting comparison between the expected population and that identified from the TESS sample is presented in Figure 7.
@@ 6.2 Approximate Bayesian Computation (ABC)
A prior probability distribution, we adopt the per-bin polynomial distribution as implemented in (Kunimoto & Matthews, 2020)
@@
A forward modeled population, used to generate a simulated set of data to compare to observations
@@
A distance function, used to determine the consensus between the simulated data and the observations
@@
We use the prior probability distribution described in Kunimoto & Matthews (2020) and modify their forward modeling method and distance function to accurately represent the characteristics of the TESS data, rather than Kepler data. Typically, in Kepler occurrence studies, the recoverability of transit signals is computed by a function dependent on their estimated signal-to-noise (SNR) based on the injection and recovery experiments (e.g. Christiansen et al., 2015). In our study, we also require recoverability to be a function of the host star spectral type, such that the differences in stellar variability and transit duration are accounted for. This approach is also computationally more efficient compared to the per-star recovery map method detailed in our section 6.1. We use one realization of ABC-PMC run to test that the SNR recovery curve approach yields similar occurrence rates compared to the per star recovery map approach and use the recovery curves for the final occurrence rate results presented in section 7.
@@
We follow Kovács et al. (2002) to derive the expected signal-to-noise of an injected transit signal for the SPOC observations. We adopt a similar signal-to-noise calculation for the QLP light curves, modified to account for its varying cadences (30- and 10-minutes),55530- and 10-minute observations, at the time we downloaded the light curves there were no 200-second FFI light curves made available on MAST we modify the calculation as follows:
@@
For each injected planet with transit depth , we split the light curves of the star into two segments with the same cadence, and estimate the number of transits, , and the average number of points in transit, for each segment. The expected SNR is then expressed as:
@@
where is the per point noise level, and is calculated as the median average deviation of the light curve.
@@
Figure 8 shows our computed recoverability as a function of SNR by stellar type in the QLP and SPOC survey. Our recoverability for both surveys plateau at .
@@
In addition, we adopt a distance function similar to Kunimoto & Matthews (2020) Equation 21 but require the computation to be done over the radius, period, and stellar type bins, rather than just radius and period. This allows us to account for the differences in stellar types in our parent stellar samples.
@@
Following Hsu et al. (2019) and Kunimoto & Matthews (2020), we first fit for the occurrence rates over a fine period-radius grid. We break the period-radius space into bins between periods of 0.78, 1.56, 3.125, 6.25, 12.5, and 20 days and radii of 0.5, 0.71, 1.41, 2.0, 2.83, 4.0, 5.66, 8.0, 11.31, and 16 R. For each period-radius cell, the occurrence rate is obtained from fitting five radius bins (two on each side in radius dimension) simultaneously. The edge radius bins ( , ) are then excluded from the reported occurrence rates. These finer resolution occurrence rate values are then combined into the final occurrence rates we report in larger cells in section 7.
@@
To account for the radius uncertainty of each planet and false positive probability of the unconfirmed planet candidates in our sample, we perform multiple realizations of the occurrence rate calculations. At each realization, the radius of each planet is drawn from a Gaussian distribution about their fitted mean and standard deviation values. Each unconfirmed planet candidate is included, or excluded, based on a random draw as per their false positive probability (see Section 4.4). Ten realizations of the SPOC and QLP samples are performed, with their resulting posteriors concatenated to attain the final occurrence rate posterior distribution, such that the uncertainties associated with planet radius and false positive rate are incorporated into our reported occurrence rates.
@@
Additionally, we account for the possibility of stars falsely identified as members of kinematically young comoving groups in our parent stellar population. Our survey uses catalogs collated from Gaia kinematic associations described in Gagné et al. (2018), Kounkel & Covey (2019), Ujjwal et al. (2020), and Moranta et al. (2022). Our analysis of the rotational variability of the stars in our parent sample indicates of the stellar population does not exhibit rotational variations consistent with the literature ages. To incorporate this uncertainty in our parent stellar population, we randomly sample of the parent stellar population in each of our ten realizations to estimate our occurrence rates. Each of the identified planet candidates and confirmed planets show signs of variability consistent with literature ages, as such we do not resample the planet population.
@@
Due to our visual vetting process, we induce non-reproducible biases, especially near the detection limit, as signals induced by smaller planets are more challenging to identify by eye. This would manifest as an under-representation of the derived occurrence rates for small planets. In this work, we do not correct our completeness maps to account for the possibility of planets being mislabeled as false-positives or false-alarms in our vetting process. We therefore do not attempt to derive an occurrence rate for the bin, which would be most heavily impacted by this bias.
6.3 Comparing occurrence rates and forward modeling expectation
@@
The forward modeled planet yield agrees well with that recovered from the young planet population for mini-Neptunes, and for short period ( day) planets. This agreement is also reflected in the occurrence rate results, where the differences between our survey results and that from the Kepler sample differ by . We find an excess of super-Neptunes, and planets at longer period orbits compared to that expected from the forward model, resulting in a noteably higher occurrence rate for these sub-populations as discussed in Section 7.1.
6.4 Comparing QLP and SPOC results
@@
The derived occurrence rates for our SPOC and QLP samples are presented in Figure 9. There exist differences between the QLP and SPOC samples at the level. We find a generally lower planet occurrence rate in the QLP sample compared to SPOC. The small planet and parent stellar population sample size makes it difficult to interpret the validity of this difference.
@@
Figure 2 shows the QLP parent sample is skewed towards earlier type stars. Kunimoto & Matthews (2020) finds K dwarfs to host more planets in the range compared to F dwarfs. The potential differences in the QLP and SPOC occurrence rates highlight the difficulty in a direct interpretation of planet frequency results without also accounting for the bulk parent stellar type. If such stellar type dependencies are validated, it points to the paucity of small planets around F stars being developed early in their evolutionary history, as expected due to the differing levels of irradiation received by these planets around early-type stars.
@@
Target stars selected as a part of the initial TESS mission and TESS Guest Investigator Programs to have their light curves sampled at 2-minute cadences may bias the derived SPOC occurrence rate. These young stars may have been known to host planets (i.e. V1298 Tau) or have been identified as planet-candidate hosts in the FFIs subsequently receiving SPOC observations during the extended TESS mission, resulting in an overall higher occurrence rate compared to the QLP sample. Further, the SPOC sample favors later-type stars, which are known to have higher occurrence rates (e.g. Kunimoto & Matthews, 2020).
@@
It is also likely the contamination rate for cluster membership of the parent population is different for QLP and SPOC. The low occurrence rate of planets in the QLP sample is more consistent with that of the mature planet population. To investigate the impact of our false positive rate for cluster membership on the derived occurrence rates, we repeat our occurrence rate calculations as described in section 6 without resampling the parent stellar population for the assumed contamination rate. Using the full stellar population observed by SPOC, we derive an occurrence rate of 28% for mini-Neptunes and 25% for super-Neptunes. For the full QLP population, we calculate an occurrence rate of for mini-Neptunes and for super-Neptunes. The excess of super-Neptunes is still prevalent in these calculations indicating that the uncertainties on our occurrence rates (see section 7.1) are derived from the small planet number statistics rather than the membership contamination rate of our parent sample. Similarly, we find the excess of Neptune-sized planets with periods of day as well in the full stellar sample. We derive an occurrence rate of 17% for the day bin with the SPOC sample and with QLP. Secure young star membership lists may help future studies resolve the discrepancies between the derived SPOC and QLP occurrence rates.
@@ 7 Occurrence rate results
The computed occurrence rates are shown in Figure 9. We breakdown the occurrence rates as a function of planet radius over the mini-Neptune () and super-Neptune () ranges. Due to only detecting two super-Earth in our SPOC and QLP surveys, we do not attempt to derive an occurrence rate for the bin. We also break down the occurrence rates as a function of orbital period, over the bins of , , , and days.
For short-period ( days), small planets ( R) younger than 200 Myr, we derive an occurrence rate of from our SPOC survey and from our QLP survey. For comparison, the Kepler occurrence rate for similar planets in this period range is . This result is qualitatively consistent with previous works that show an increase in the planet occurrence rate about young stars (e.g. Rizzuto et al., 2017; Fernandes et al., 2023; Christiansen et al., 2023).
@@ 7.1 The planet distribution about young stars is different to that of the mature population
In both the QLP and SPOC samples, we find (1) a mild excess of super-Neptunes in the young planet population, and (2) a significant excess of Neptune-sized planets at day orbital periods in the young planet population.
@@ 7.1.1 We find a mild excess of super-Neptunes orbiting young stars
Previous young planet censuses have found a higher occurrence rate of super-Neptunes (e.g. Rizzuto et al., 2017; Fernandes et al., 2023; Christiansen et al., 2023) at ages up to that of the Hyades cluster. David et al. (2021) and Berger et al. (2023) found the distribution of small planets is continuously modified over the billion-year timescale, pointing to the role core-powered mechanisms play in shaping the mature planet population.
Our work presents the radius distribuiton of planets amongst one of the youngest stellar distributions studied so far. We find a mild excess of super-Neptune-sized planets around the youngest surveyed population in literature. The Kepler sample yielded an occurrence rate of for super-Neptunes in the mature population (Kunimoto & Matthews, 2020). We derive higher occurrence rates for super-Neptunes (SPOC differing to the equivalent Kepler occurrence rate at significance; QLP % differing from Kepler at ) in our Myr sample (see Figure 9). A similar significance exists in the forward modeling comparison, at for SPOC and for QLP. These qualitatively agree with previous works showing an excess of inflated planets around young stars.
@@ 7.1.2 We find a significant excess of day period planets around young stars
The period distribution of young planets appears distinctively different from that of the mature population. Within our SPOC sample, we predict 0.94 planets between days and find 7. The same excess of day period planets is also apparent in our QLP sample. The Kepler statistic predicts 1.6 planets between days, while we recover 9 such planets.
@@
This apparent period evolution between our million year sample and that of the Kepler distribution helps us understand the mechanisms that shape the sub-Saturn desert. Hallatt & Lee (2022) and Thorngren et al. (2023) find that XUV-driven mass-loss can lead to runaway envelope stripping of the least dense hot Saturns. In this scenario, close-in planets with initial radii undergo runaway mass loss within 100 Myr. In addition, close-in planets that have undergone runaway mass-loss remain puffy at the radius due to residual heat, before cooling and contracting to at time scales of over Myr.
@@
Hansen & Murray (2013) finds that an initial over-density of planets in the day orbital period range is a natural result of in-situ formation of rocky cores with subsequent accretion of gaseous envelopes. This distance corresponds to the dust sublimation radius for Sun-like stars (Swift et al., 2013). Tidal evolution models from Hansen & Murray (2015) show that this initial buildup of planets at days can migrate inward via tidal dissipation and eccentricity damping at the billion-year timescale if they exist in compact multiplanet systems that mutually undergo eccentricity excitation. Such tidally driven effects may be responsible for smoothing the over-density of young planets into the mature planet period distribution.
7.2 Comparison to Hydrogen-helium dominated planet evolution models

### 11 Apr 2024 v2

@@ 6.1 Forward Modeling
We perform a forward modeling test to compare the demographics of planets around young stars against that of the Kepler distribution. We synthesize a population of planets using the Kepler statistics as presented in Kunimoto & Matthews (2020) and compute the expected number of recovered planets as per our survey efficiency. The forward modeling process allows us to compare the expected characteristics of the planet population against those that we identify from the young stellar sample.
We adopt the Kunimoto & Matthews (2020) Kepler occurrence rates for F, G, and K stars. We draw the occurrence rates assuming half-Gaussians based on the asymmetric uncertainties in Kunimoto & Matthews (2020) for each of the simulations. For each star in our parent sample, we simulate a planetary system based on its Kepler period-radius planet frequency grid for its specific stellar type.444The definition of the stellar-type boundaries are adopted to be the same as Kunimoto & Matthews (2020) For the M dwarfs in our sample, we use the occurrence rates for K stars to approximate their planet occurrence rate.. The planets are first randomly assigned into a period-radius cell based on the two-dimensional cumulative probability. Then, their period and radius are drawn within the cell assuming log-uniform distributions in both axes. We then account for transit probability and detection probability as per the injection and recovery exercise described in Section 5. For each planet, the transit probability is calculated as , as we reject near-grazing candidates with impact parameters of due to their difficulty for confirmation. Detection probability is computed as the percentage of recovered planets within a single period-radius bin as shown in Figure 6 from our injection and recovery exercise. This forward modeling process is performed 100 times each for the QLP and SPOC samples to quantify its associated uncertainties. The resulting comparison between the expected population and that identified from the TESS sample is presented in Figure 7.
@@ 6.2 Approximate Bayesian Computation (ABC)
a prior probability distribution, where we adopt the per-bin polynomial distribution as implemented in Kunimoto & Matthews (2020),
@@
a forward modeled population, used to generate a simulated set of data to compare to observations, and
@@
a distance function, used to determine the consensus between the simulated data and the observations.
@@
We use the prior probability distribution described in Kunimoto & Matthews (2020) and modify their forward modeling method and distance function to accurately represent the characteristics of the TESS data, rather than Kepler data. Typically, in Kepler occurrence studies, the recoverability of transit signals is computed by a function dependent on their estimated S/Ns based on the injection and recovery experiments (e.g. Christiansen et al., 2015). In our study, we also require recoverability to be a function of the host star spectral type, such that the differences in stellar variability and transit duration are accounted for. This approach is also computationally more efficient compared to the per-star recovery map method detailed in our section 6.1. We use one realization of an ABC-PMC run to test that the S/N recovery curve approach yields similar occurrence rates compared to the per-star recovery map approach and use the recovery curves for the final occurrence rate results presented in section 7.
@@
We follow Kovács et al. (2002) to derive the expected S/N of an injected transit signal for the SPOC observations. We adopt a similar S/N calculation for the QLP light curves, modified to account for its varying cadences (30 and 10 minutes),55530 and 10 minute observations; at the time we downloaded the light curves there were no 200 s FFI light curves made available on MAST and we modify the calculation as follows.
@@
For each injected planet with transit depth , we split the light curves of the star into two segments with the same cadence, and estimate the number of transits, , and the average number of points in transit, for each segment. The expected S/N is then expressed as:
@@
where is the per-point noise level, and is calculated as the median average deviation of the light curve.
@@
Figure 8 shows our computed recoverability as a function of S/N by stellar type in the QLP and SPOC survey. Our recoverability for both surveys plateaus at .
@@
In addition, we adopt a distance function similar to Kunimoto & Matthews (2020) Equation 21 but require the computation to be done over the radius, period, and stellar-type bins, rather than just radius and period. This allows us to account for the differences in stellar types in our parent stellar samples.
@@
Following Hsu et al. (2019) and Kunimoto & Matthews (2020), we first fit for the occurrence rates over a fine period-radius grid. We break the period-radius space into bins between periods of 0.78, 1.56, 3.125, 6.25, 12.5, and 20 days and radii of 0.5, 0.71, 1.41, 2.0, 2.83, 4.0, 5.66, 8.0, 11.31, and 16 . For each period-radius cell, the occurrence rate is obtained from fitting five radius bins (two on each side in radius dimension) simultaneously. The edge radius bins (, ) are then excluded from the reported occurrence rates. These finer-resolution occurrence rate values are then combined into the final occurrence rates we report in larger cells in section 7.
@@
To account for the radius uncertainty of each planet and false-positive probability of the unconfirmed planet candidates in our sample, we perform multiple realizations of the occurrence rate calculations. At each realization, the radius of each planet is drawn from a Gaussian distribution about their fitted mean and standard deviation values. Each unconfirmed planet candidate is included, or excluded, based on a random draw as per their false-positive probability (see Section 4.4). 10 realizations of the SPOC and QLP samples are performed, with their resulting posteriors concatenated to attain the final occurrence rate posterior distribution, such that the uncertainties associated with planet radius and false positive rate are incorporated into our reported occurrence rates.
@@
Additionally, we account for the possibility of stars falsely identified as members of kinematically young comoving groups in our parent stellar population. Our survey uses catalogs collated from Gaia kinematic associations described in Gagné et al. (2018), Kounkel & Covey (2019), Ujjwal et al. (2020), and Moranta et al. (2022). Our analysis of the rotational variability of the stars in our parent sample indicates of the stellar population does not exhibit rotational variations consistent with the literature ages. To incorporate this uncertainty in our parent stellar population, we randomly sample of the parent stellar population in each of our 10 realizations to estimate our occurrence rates. Each of the identified planet candidates and confirmed planets show signs of variability consistent with literature ages; as such we do not resample the planet population.
@@
Due to our visual vetting process, we induce nonreproducible biases, especially near the detection limit, as signals induced by smaller planets are more challenging to identify by eye. This would manifest as an underrepresentation of the derived occurrence rates for small planets. In this work, we do not correct our completeness maps to account for the possibility of planets being mislabeled as false positives or false alarms in our vetting process. We therefore do not attempt to derive an occurrence rate for the bin, which would be most heavily impacted by this bias.
6.3 Comparing Occurrence Rates and Forward Modeling Expectations
@@
The forward modeled planet yield agrees well with that recovered from the young planet population for mini-Neptunes, and for short-period ( day) planets. This agreement is also reflected in the occurrence rate results, where the differences between our survey results and those from the Kepler sample differ by . We find an excess of super-Neptunes, and planets at longer period orbits compared to that expected from the forward model, resulting in a noteably higher occurrence rate for these subpopulations as discussed in Section 7.1.
6.4 Comparing the QLP and SPOC Results
@@
The derived occurrence rates for our SPOC and QLP samples are presented in Figure 9. There exist differences between the QLP and SPOC samples at the level. We find a generally lower planet occurrence rate in the QLP sample compared to the SPOC sample. The small planet and parent stellar population sample size makes it difficult to interpret the validity of this difference.
@@
Figure 2 shows the QLP parent sample is skewed toward earlier-type stars. Kunimoto & Matthews (2020) find K dwarfs to host more planets in the range compared to F dwarfs. The potential differences in the QLP and SPOC occurrence rates highlight the difficulty in a direct interpretation of planet frequency results without also accounting for the bulk parent stellar type. If such stellar-type dependencies are validated, it points to the paucity of small planets around F stars being developed early in their evolutionary history, as expected due to the differing levels of irradiation received by these planets around early-type stars.
@@
Target stars selected as a part of the initial TESS mission and TESS Guest Investigator Programs to have their light curves sampled at 2 minute cadences may bias the derived SPOC occurrence rate. These young stars may have been known to host planets (i.e., V1298 Tau) or have been identified as planet candidate hosts in the FFIs subsequently receiving SPOC observations during the extended TESS mission, resulting in an overall higher occurrence rate compared to the QLP sample. Further, the SPOC sample favors later-type stars, which are known to have higher occurrence rates (e.g. Kunimoto & Matthews, 2020).
@@
It is also likely the contamination rate for cluster membership of the parent population is different for QLP and SPOC. The low occurrence rate of planets in the QLP sample is more consistent with that of the mature planet population. To investigate the impact of our false-positive rate for cluster membership on the derived occurrence rates, we repeat our occurrence rate calculations as described in section 6 without resampling the parent stellar population for the assumed contamination rate. Using the full stellar population observed by SPOC, we derive an occurrence rate of 28% for mini-Neptunes and 25% for super-Neptunes. For the full QLP population, we calculate an occurrence rate of for mini-Neptunes and for super-Neptunes. The excess of super-Neptunes is still prevalent in these calculations; indicating that the uncertainties on our occurrence rates (see section 7.1) are derived from the small planet number statistics rather than the membership contamination rate of our parent sample. Similarly, we find the excess of Neptune-sized planets with periods of day as well in the full stellar sample. We derive an occurrence rate of 17% for the day bin with the SPOC sample and with QLP. Secure young star membership lists may help future studies resolve the discrepancies between the derived SPOC and QLP occurrence rates.
@@ 7 Occurrence rate results
The computed occurrence rates are shown in Figure 9. We break down the occurrence rates as a function of planet radius over the mini-Neptune () and super-Neptune () ranges. Due to only detecting two super-Earths in our SPOC and QLP surveys, we do not attempt to derive an occurrence rate for the bin. We also break down the occurrence rates as a function of orbital period, over the bins of , , , and days.
For short-period ( days), small planets () younger than 200 Myr, we derive an occurrence rate of from our SPOC survey and from our QLP survey. For comparison, the Kepler occurrence rate for similar planets in this period range is . This result is qualitatively consistent with previous works that show an increase in the planet occurrence rate about young stars (e.g. Rizzuto et al., 2017; Christiansen et al., 2023; Fernandes et al., 2023).
@@ 7.1 The Planet Distribution about Young Stars is Different to that of the Mature Population
In both the QLP and SPOC samples, we find (1) a mild excess of super-Neptunes in the young planet population and (2) a significant excess of Neptune-sized planets at day orbital periods in the young planet population.
@@ 7.1.1 We Find a Mild Excess of Super-Neptunes Orbiting Young Stars
Previous young planet censuses have found a higher occurrence rate of super-Neptunes (e.g. Rizzuto et al., 2017; Christiansen et al., 2023; Fernandes et al., 2023) at ages up to that of the Hyades cluster. David et al. (2021) and Berger et al. (2023) found the distribution of small planets is continuously modified over a billion-year timescale, pointing to the role core-powered mechanisms play in shaping the mature planet population.
Our work presents the radius distribution of planets amongst one of the youngest stellar distributions studied so far. We find a mild excess of super-Neptune-sized planets around the youngest surveyed population in the literature. The Kepler sample yielded an occurrence rate of for super-Neptunes in the mature population (Kunimoto & Matthews, 2020). We derive higher occurrence rates for super-Neptunes (SPOC: , differing to the equivalent Kepler occurrence rate at significance; QLP: %, differing from Kepler at ) in our Myr sample (see Figure 9). A similar significance exists in the forward modeling comparison, at for SPOC and for QLP. These qualitatively agree with previous works showing an excess of inflated planets around young stars.
@@ 7.1.2 We Find a Significant Excess of day Period Planets Around Young Stars
The period distribution of young planets appears distinctively different from that of the mature population. Within our SPOC sample, we predict 0.94 planets between 6.2 and 12 days and find seven. The same excess of day period planets is also apparent in our QLP sample. The Kepler statistic predicts 1.6 planets between 6.2 and 12 days, while we recover nine such planets.
@@
This apparent period evolution between our Myr sample and that of the Kepler distribution helps us understand the mechanisms that shape the sub-Saturn desert. Hallatt & Lee (2022) and Thorngren et al. (2023) find that XUV-driven mass loss can lead to runaway envelope stripping of the least dense hot Saturns. In this scenario, close-in planets with initial radii undergo runaway mass loss within 100 Myr. In addition, close-in planets that have undergone runaway mass loss remain puffy at the radius due to residual heat, before cooling and contracting to at time scales of over Myr.
@@
Hansen & Murray (2013) find that an initial overdensity of planets in the day orbital period range is a natural result of the in-situ formation of rocky cores with the subsequent accretion of gaseous envelopes. This distance corresponds to the dust sublimation radius for Sun-like stars (Swift et al., 2013). Tidal evolution models from Hansen & Murray (2015) show that this initial buildup of planets at days can migrate inward via tidal dissipation and eccentricity damping over a billion-year timescale if they exist in compact multiplanet systems that mutually undergo eccentricity excitation. Such tidally driven effects may be responsible for smoothing the overdensity of young planets into the mature planet period distribution.
7.2 Comparison to Hydrogen-Helium-dominated Planet Evolution Models